OpenStudy (anonymous):

i vneed help finding the y intercept of this equation. 5x-8y=120

mathslover (mathslover):

note that when we are given to find y intercept.. we have to put x = 0 and then solve for y

OpenStudy (anonymous):

y-intercept is when x=0 so substitute x=0 into the equation to find y.

mathslover (mathslover):

so we have : 5x-8y = 120 5(0) - 8y=120 -8y = 120

mathslover (mathslover):

can you solve for y ? @Renmack

OpenStudy (anonymous):

yep, thanks for the help...now that I read the answer I realise I know this - brain is obviously not working.